OpenStudy (anonymous):

Find the slope by solving for y -3x + 2y = 9 show all steps

OpenStudy (anonymous):

-3x + 2y = 9 +3x +3x 2y = 3x + 9 /2 /2 y = 3/2x + 9/2

OpenStudy (anonymous):

YAY I GOT IT RIGHT!! :)

OpenStudy (anonymous):

\[ \frac{3}{2} \]